OpenStudy (anonymous):

how would i condense Log8 12 + Log 8 X

OpenStudy (callisto):

Log(8*12) + Log (8X) ?

OpenStudy (campbell_st):

log laws Log(ab) = log(a) + log(b) so do the reverse \[\log _{8} (12x) = \log _{8}(12) + \log _{8} (x)\]

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

okay i got that now but what about this one Log3(x^4y^6/4) that would be log3(x^4/4)-log3(y^6/4)

OpenStudy (campbell_st):

log(a) - log(b) = log(a/b) so its \[\log _{3} (x^4y^6) - \log _{3}(4)\]

OpenStudy (campbell_st):

then it becomes \[\log _{3}(x^4) + \log _{3}(y^6) - \log _{3}(4)\]

OpenStudy (anonymous):

woah

OpenStudy (anonymous):

so you compleatly split it up?

OpenStudy (campbell_st):

but then log (a^x) = xlog(a) so you question could be \[4\log _{3}(x) + 6\log _{3}(y) - \log _{3}(4)\]

OpenStudy (anonymous):

so those would be the quotient property then the power property? right?

OpenStudy (campbell_st):

quotient. product and power

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

yet why is it a negative?